OpenStudy (anonymous):

Ok so I am graphing and need to find the two plot points for f(x)=4x-6 Can anyone walk me through finding this I beleve you can substitute any number for X but my math is a mess and I cant figure it out

OpenStudy (anonymous):

let y=f(x)

OpenStudy (anonymous):

y=4x-6

OpenStudy (anonymous):

set x equals to 0 y=4(0)-6 y=-6 (0,-6)

OpenStudy (anonymous):

yes then put in for x right such as 0-4 is it

OpenStudy (anonymous):

3/2 is not right on the graph 0-6 is how do I find the second point to plot i understand yur math but cant figure out the secound number

OpenStudy (anonymous):

how do u figure that out when you need two numbers 0,-6 was a point but to find the other

OpenStudy (anonymous):

I need a tutor im confused ugh

OpenStudy (anonymous):

set y=-6 x=0 (0,-6)